THE STRUCTURE OF CO_{2}-N_{2}O DETERMINED BY FOURIER TRANSFORM MICROWAVE SPEC-TROSCOPY OF THE ^{15}N^{14}NO-CONTAINING ISOTOPOMER

\underline{MARK D. MARSHALL}, Department of Chemistry-2243, Amherst College, P.O. Box 5000, Amherst, MA 01002-5000; and HELEN O. LEUNG, Department of Chemistry, Mount Holyoke College, South Hadley, MA 01075-6407.

In both the infrared ^{a} and previous microwave ^{b} studies of the van der Waals molecule CO_{2} − N_{2}O, it was impossible to resolve experimentally the structural ambiguity concerning the orientation of the N_{2}O subunit in this approximately slipped parallel complex. Isotopic substitution of the outer nitrogen in the N_{2}O results in a microwave spectrum that is consistent with a structure in which the terminal oxygen of the N_{2}O is closer to the central carbon of the CO_{2}. Substitution coordinates derived from Kraitchman’s equations support this conclusion. This experimentally determined structure is in agreement with the conclusion reached on the basis of ab initio results used in conjunction with the earlier infrared work.^{a}

aC. Dutton, A. Sazonov, and R. A. Beaudet, J. Phys. Chem. \underline{100,} 17772 (1996).

bH. O. Leung, J. Chem. Phys. \underline{108,} xxxx (1998).